# Graph Linearization Methods for Reasoning on Graphs with Large Language Models

### v1

@@ Appendix A Prompt templates
In an undirected graph, (i, j) means that node i and node j are connected with an undirected edge. The edges in G are: {linearized graph}
Q: How many nodes are in this graph?
Max Degree Calculation:
@@
In an undirected graph, (i, j) means that node i and node j are connected with an undirected edge. The degree of a node is the number of edges connected to the node. Given a graph G and it’s list of edges respond to the following question:
@@
Q: Without any justification, what is the maximum node degree in the following graph G? G: {linearized graph}
@@
Q: Which of the defined structures is included in the following graph? graph: {linearized graph}

### v3

@@ Appendix D Prompt templates
In an undirected graph $G$, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge.
Q: How many nodes are in $G$?
G: {linearized graph}
Max Degree
@@
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. The degree of a node is the number of edges connected to the node. Given a graph $G$ and its list of edges, respond to the following question:
@@
Q: Without any justification, what is the maximum node degree in the following graph $G$?
G: {linearized graph}
Node Degree
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. The degree of a node is the number of edges connected to the node. Given a graph $G$ and its list of edges, respond to the following question:
Q: Without any justification, what is the degree of node {node} in the following graph $G$?
G: {linearized graph}
Edge Existence
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. Given a graph $G$ and its list of edges, respond to the following question:
Q: Does an undirected edge ($(\{node1\},\{node2\})$ exist in the following graph $G$?.
G: {linearized graph}
Diameter
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. The diameter of a graph is the length of the shortest path between the most distanced nodes. Given a graph $G$ and its list of edges, respond to the following question:
Q: Without any justification, what is the diameter of the following graph $G$?
G: {linearized graph}
Shortest Path
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. Given a graph $G$ and its list of edges, respond to the following question:
Q: Without any justification, what is the length of the shortest path from node $\{node1\}$ to node $\{node2\}$? If no path exists, the response is ’0’.
G: {linearized graph}
Path Existence
In an undirected graph, $(i,j)$ means that node $i$ and node $j$ are connected with an undirected edge. Given a graph $G$ and its list of edges, respond to the following question:
Q: Does a path that connects node $\{node1\}$ and $\{node2\}$ exist in the following graph $G$?
G: {linearized graph}
@@
Q: Which of the defined structures is included in the following graph?
graph: {linearized graph}